OpenStudy (anonymous):

Solve: 5(x − 1) − (x + 5) = 2 1/2 3 2 1

OpenStudy (nubeer):

5x-5-x-5=2 distributive property.. now rest is simple algebra ..

OpenStudy (mew55):

like nubeer said, distribute first so 5(x-1)-(x+5)=2 *keep in mind, the minus sign changes the signs when distributing so 5(x-1)-(x+5)=2 5x-5-x-5=2 2. collect like terms 4x *the -5-5 changes to an addition problem cause of the negative -5-5=-5+-5=-10 4x-10=2 3.add 10 to both sides 4x=12 4.divide both sides x=3